# Arrow-Debreu Meets Kyle: Price Discovery Across Derivatives

### v28

@@ 1 Introduction
Across option markets, trades in one option often impact the prices of others as options trades jointly convey information about the underlying payoff. For example, a common volatility trade involves trading a put-call pair to profit from price movements regardless of direction. Therefore, trades in such pairs jointly convey volatility information. Our characterization of such cross-market information spillovers generates new empirical predictions linking cross-option dynamics to various aspects of the underlying payoff. Moreover, we offer an equilibrium explanation for the “volatility smile,” an empirical regularity of option-implied volatilities that has been widely studied but only in reduced-form.
@@
Many option pricing and financial econometric models have sought to capture the volatility smile by correcting misspecifications in the Black and Scholes (1973) framework. Prominent classes of such models include the jump-diffusion, [^1] See Merton (1976) and Bates (1991). stochastic volatility, [^2] See Heston (1993), Bates (1996), Duffie et al. (2000), Britten-Jones and Neuberger (2000), Bates (2000), Aït-Sahalia and Kimmel (2007), Christoffersen et al. (2010), and Aït-Sahalia et al. (2021). and local volatility models. [^3] See Dupire (1994), Berestycki et al. (2002), and Carr and Cousot (2012). Our framework moves beyond reduced-form model misspecifications to directly explain the strike structure of implied volatility via cross-strike information dynamics.
The rest of the paper proceeds as follows. Section 2 demonstrates the basic intuition of our results through a simplified setting. Section 3 presents the general model. Section 4 identifies the sufficient statistic for the uninformed trader’s cross-market inferences. Section 5 analyzes the informed trader’s cross-market price impact. Section 6 establishes equilibrium through a canonical reduction of the general trading game. While this canonical reduction is a key modeling innovation that makes our general model tractable, the broad intuition of the resulting equilibrium extends directly from Section 2 and can be understood from Example 6.7 without delving into the precise details. Section 7 discusses the characterization of price discovery across contingent claims—the informed demand, price impact, and information efficiency of prices. Section 8 explores the immediate empirical implications of our results for price discovery across option markets. Section 9 concludes. Formal assumptions and proofs are provided in the Appendix.
@@
Insider Expected Profit Curve. The insider’s optimal portfolio size $\alpha^{*}$ characterizes equilibrium.
@@ 2 Basic Intuition
The insider’s optimal portfolio size $\alpha^{*}$ trades off the marginal payoff from exploiting the true distribution $(\frac{1}{2},0,\frac{1}{2})$ against the marginal costs. These costs include the cost incurred at current prices, assuming no price impact, and the cost stemming from the cross-market price impact due to information revelation. His optimal portfolio size $\alpha^{*}$ characterizes equilibrium—this simple result will extend to the general model. Allowing correlated noise trades, varying noise intensity, and non-uniform priors does not materially affect the results. [^6] Noise Trades Correlation Allowing noise trades to be correlated, while keeping their intensity constant, does not affect price impact. For example, suppose the correlation of noise trades for a pair of assets has an opposite sign to their payoff correlation. In that case, the insider would scale up his trading strategy to capture additional losses incurred by the noise traders trading against the payoffs. The market maker would shift his supply schedules accordingly, but the price impact at the margin would remain the same; this is similar to biased noise trades in the single-asset setting.  [^7] Noise Trading Intensity Suppose the noise trading intensity $(\sigma_{i})$ varies across markets, with $\epsilon_{i}\!\stackrel{{\scriptstyle d}}{{\sim}}\!\mathcal{N}(0,\sigma_{i}^{2})$, $i=1,2,3$. Then $\Delta(\omega)$ becomes $\Delta(\omega)=2(\frac{\omega_{1}}{\sigma_{1}^{2}}+\frac{\omega_{3}}{\sigma_{3}^{2}}-\frac{\omega_{2}}{\sigma_{2}^{2}}).$ Thus, as expected, higher noise trading intensity reduces price impact within each market and across markets.  [^8] General Prior When the market maker has a general prior $(\pi_{0}(s_{1}),\pi_{0}(s_{2}))$, the posterior likelihood ratio (1) becomes $\frac{\pi_{1}(s_{1}|\omega)\pi_{0}(s_{1})}{\pi_{1}(s_{2}|\omega)\pi_{0}(s_{2})}=\frac{e^{\alpha\omega_{1}-\frac{1}{2}\alpha^{2}-\alpha\omega_{2}-\frac{1}{2}\alpha^{2}+\alpha\omega_{3}-\frac{1}{2}\alpha^{2}}\pi_{0}(s_{1})}{e^{-\alpha\omega_{1}-\frac{1}{2}\alpha^{2}+\alpha\omega_{2}-\frac{1}{2}\alpha^{2}-\alpha\omega_{3}-\frac{1}{2}\alpha^{2}}\pi_{0}(s_{2})}.$ The comparative statics for the prior follows intuitively. For example, a higher $\pi_{0}(s_{1})$ leads to greater cross price impact between securities $x_{1}$ and $x_{3}$—because it leads the market maker to place more emphasis on the co-movement of the orders for $x_{1}$ and $x_{3}$ in his inference.
@@ 4.2 Sufficient Statistics
The market maker adjusts for the noise trading intensities $(\sigma_{i})$ in making cross-market inferences. Higher noise levels reduce the informativeness of his sufficient statistic $(\sum\limits_{i}\frac{\widetilde{W}(x_{i}|s_{k})\omega_{i}}{\sigma_{i}^{2}})_{s_{k}\in S}$. In the limit as $\sigma_{k}\!\rightarrow\!\infty$ for each $k$, this statistic reduces to the trivial zero statistic, and his posterior (11) reduces to the prior $\pi_{0}$, i.e., he infers no information from order flow.
@@ 5.3 No-Arbitrage
No-arbitrage is a general necessary condition for the existence of equilibrium. We now state this condition for our setting. Here, once again, the projection coefficient profile $\Pi_{insider}$ between the insider’s trades and the market maker’s belief naturally occurs. Given a market maker belief $\widetilde{W}$, we define the corresponding zero price impact portfolios as portfolios $W$ for which $\Pi_{insider}(W,\,\cdot\,;\widetilde{W})=0$. Intuitively, such a portfolio causes zero price impact because it is orthogonal to the market maker’s belief about the insider’s trading strategy. Such a portfolio must yield a zero payoff for the insider; otherwise, arbitrage opportunities arise for the insider. This is Theorem 5.4.
@@
Given any market maker belief, an optimal portfolio for the insider can exist only if the zero price impact portfolios corresponding to that belief yield zero payoff for the insider. Otherwise, the insider can obtain unbounded utility.
@@ 6.1 Canonical Game
The zero payoff portfolios are those orthogonal to the linear span of $\{\eta(\,\cdot\,,s_{k})\}_{k}$—any portfolio $W$ satisfying $\sum\limits_{i}W(x_{i})\eta(x_{i}|s_{k})=0$ for all $s_{k}$ yields zero payoff for the insider. Thus, the insider’s payoff-relevant portfolios lie within the linear span of $\{\eta(\,\cdot\,|s_{k})\}_{k}$. We state this result as Proposition 6.1(i). On the other hand, the market maker’s noise-adjusted equilibrium beliefs must contain this span; otherwise, zero price impact allows arbitrage by the insider. This is Proposition 6.1(ii).
@@ 7.1 Informed Demand
informed $\text{{\em informed demand}}\,\begin{cases}\text{{\em single-asset} :}&\frac{\sigma_{\varepsilon}}{\sigma_{v}}\cdot(s-v_{0})\\[2.58334pt] \text{{\em contingent claims} :}&{\bf L}^{-1}\cdot\alpha^{*}{\bf Q}\end{cases}.$
  • Information Intensity ($\frac{\sigma_{\varepsilon}}{\sigma_{v}}$ vs. ${\bf L}$): Both $\frac{\sigma_{v}}{\sigma_{\varepsilon}}$ and ${\bf L}$ quantify the insider’s information intensity in their respective settings. In the single-asset setting, a larger $\frac{\sigma_{v}}{\sigma_{\varepsilon}}$ means the market maker can more effectively infer the variation in the asset payoff from the variation in order flow, thus reducing informed demand. For contingent claims, a larger ${\bf L}$ means the market maker can better infer the variations in the payoff distributions across signals from the variations in cross-market order flow, which reduces informed demand.
@@
The following examples illustrate the informed demand for trading on the first, second, and third moments of an underlying asset payoff. These examples explain how the corresponding option strategies used in trading practice emerge as equilibrium strategies. Other common option strategies are explained similarly under appropriate specifications. In Examples 7.1, 7.2, and 7.3, the signal is binary, i.e., $S=\{s_{1},s_{2}\}$, and all payoff distributions are suitably discretized.
@@
(a)/(b): high expected payoff signal/bull spread, (c)/(d): low expected payoff signal/bear spread.
@@
(a)/(b): high volatility signal/straddle, (c)/(d): low volatility signal/butterfly.
Price Impact Three states are indicated—$x_{1}$, $x_{2}$, and $x_{3}$.
$\bullet\,x_{2}$: zero price impact on all securities because its payoff is constant (has zero variation) across signals. Thus, the informed demand for $x_{2}$ must be zero—see (b) and (d)
@@
The left column of Figure 4 shows the possible payoff distributions, with the insider’s private signal shaded. The right column of Figure 4 shows the informed demand corresponding to each signal. The informed demands align with common option strategies for volatility trading (see Hull (2015)). When the insider anticipates high volatility, his option strategy in Figure 4b resembles a straddle or long iron condor pattern. When he anticipates low volatility, his option strategy in Figure 4d follows a butterfly pattern.
@@ 7.2 Price Impact
    Informed Demand Scale $(\alpha^{*}$): As the informed demand (33) scales up by $\alpha^{*}$, price impact responds one-to-one.
@@
    Noise Intensity Adjustment $(\frac{1}{\sqrt{\sigma_{i}\sigma_{j}}})$: Higher noise trading intensity in either market reduces cross price impact.
@@
    Payoff Covariance $(\mathbb{E}\left[{\operatorname{\mathbb{Cov}}}\left(\eta(x_{i}|\,\cdot\,),\eta(x_{j}|\,\cdot\,)\big{|}\omega\right)\right])$: The (expected) covariance of securities $x_{i}$ and $x_{j}$ payoffs under the market maker’s equilibrium posterior.
@@ 9 Conclusion
By unifying the elements of Arrow and Debreu (1954) and Kyle (1985), we have developed a general model of informed trading across contingent claim markets. Our findings advance the theoretical understanding of how contingent claim markets jointly incorporate information. Our framework provides explanations for observed empirical practices, bridges the gap with empirical literature, and proposes new avenues for future research.
@@ A.1 Model Assumptions
In other words, $\mathbb{P}_{\!\scriptscriptstyle 0}$ specifies the distribution of noise-only order flow $\omega$ when insider demand is zero across markets.
@@
The marginal utility functional $\partial\!\kern 0.80002ptJ(\,\cdot\,;W)\colon\mathbb{R}^{\infty}\rightarrow\mathbb{R}$ is the Gâteaux-derivative of the expected utility functional $J(\,\cdot\,|s_{k};\widetilde{W})\colon\mathbb{R}^{\infty}\rightarrow{\mathbb{R}}$, as defined in Equation (14), at $W$. (Both $J(\,\cdot\,|s_{k};\widetilde{W})$ and its Gâteaux-derivative $\partial\!\kern 0.80002ptJ(\,\cdot\,;W)$ are conditional on $s_{k}$ and parameterized by $\widetilde{W}$.) $J(W|s_{k};\widetilde{W})=J_{p}(W)-J_{c}(W)$ is the difference between
[…]
and
[…]

### v29

@@ 1 Introduction
Across option markets, trades in one option often impact the prices of others as options trades jointly convey information about the underlying payoff. For example, a common volatility trade involves trading a put-call pair to profit from price movements regardless of direction. Therefore, trades in such pairs jointly convey volatility information. Our characterization of such cross-market information spillovers generates new empirical predictions linking cross-option dynamics to various aspects of the underlying payoff. Moreover, we offer the first equilibrium explanation for the “volatility smile,” an empirical regularity of option-implied volatilities that has been widely studied but only in reduced-form.
@@
Many option pricing and financial econometric models have sought to capture the volatility smile by correcting various misspecifications in the Black and Scholes (1973) framework. Prominent approaches include the jump-diffusion models [^1] See Merton (1976) and Bates (1991)., stochastic volatility models [^2] See Heston (1993), Bates (1996), Duffie et al. (2000), Britten-Jones and Neuberger (2000), Bates (2000), Aït-Sahalia and Kimmel (2007), Christoffersen et al. (2010), and Aït-Sahalia et al. (2021)., and local volatility models. [^3] See Dupire (1994), Berestycki et al. (2002), and Carr and Cousot (2012). Our framework moves beyond reduced-form model misspecifications to directly explain the strike structure of implied volatility via cross-strike information dynamics.
The rest of the paper proceeds as follows. Section 2 demonstrates the basic intuition of our results through a simplified setting. Section 3 presents the general model. Section 4 identifies the intuitive sufficient statistic for the uninformed trader’s cross-market inferences. Section 5 analyzes the informed trader’s cross-market price impact. Section 6 establishes equilibrium through a canonical reduction of the general trading game. While this canonical reduction is a key modeling innovation that makes our general model tractable, the broad intuition of the resulting equilibrium extends directly from Section 2 and can be understood from Example 6.7 without delving into the precise details. Section 7 discusses the characterization of price discovery across contingent claims—the informed demand, price impact, and information efficiency of prices. Section 8 explores the immediate empirical implications of our results for price discovery across option markets. Section 9 concludes. Formal assumptions and proofs are provided in the Appendix.
@@
Insider Expected Profit Curve. The insider’s optimal portfolio size $\alpha^{*}$ is the endogenous constant characterizing equilibrium.
@@ 2 Basic Intuition
The insider’s optimal portfolio size $\alpha^{*}$ balances the marginal payoff from exploiting the true payoff distribution against the marginal cost of his trades. The marginal cost comprises that incurred at current prices, assuming no price impact, and that resulting from the cross-market price impact due to information revelation. His optimal portfolio size $\alpha^{*}$ is the endogenous constant characterizing equilibrium. This simple equilibrium structure extends to the general model (see Example 6.7). Allowing correlated noise trades, varying noise intensity, and non-uniform priors does not materially affect the results. [^6] Noise Trades Correlation Allowing noise trades to be correlated, while keeping their intensity constant, does not affect price impact. For example, suppose the correlation of noise trades for a pair of assets has an opposite sign to their payoff correlation. In that case, the insider would scale up his trading strategy to exploit the noise traders’ additional losses from trading against the payoffs. The market maker would shift his supply schedules accordingly, but the price impact at the margin remains the same—akin to having biased noise trades in the single-asset setting.  [^7] Noise Trading Intensity Suppose the noise trading intensity $(\sigma_{i})$ varies across markets, with $\epsilon_{i}\!\stackrel{{\scriptstyle d}}{{\sim}}\!\mathcal{N}(0,\sigma_{i}^{2})$, $i=1,2,3$. Then $\Delta(\omega)$ becomes $\Delta(\omega)=2(\frac{\omega_{1}}{\sigma_{1}^{2}}+\frac{\omega_{3}}{\sigma_{3}^{2}}-\frac{\omega_{2}}{\sigma_{2}^{2}}).$ Thus, as expected, higher noise trading intensity reduces price impact both within each and across markets.  [^8] General Prior When the market maker has a general prior $(\pi_{0}(s_{1}),\pi_{0}(s_{2}))$, the posterior likelihood ratio (1) becomes $\frac{\pi_{1}(s_{1}|\omega)\pi_{0}(s_{1})}{\pi_{1}(s_{2}|\omega)\pi_{0}(s_{2})}=\frac{{\rm e}^{\alpha\omega_{1}-\frac{1}{2}\alpha^{2}-\alpha\omega_{2}-\frac{1}{2}\alpha^{2}+\alpha\omega_{3}-\frac{1}{2}\alpha^{2}}\pi_{0}(s_{1})}{{\rm e}^{-\alpha\omega_{1}-\frac{1}{2}\alpha^{2}+\alpha\omega_{2}-\frac{1}{2}\alpha^{2}-\alpha\omega_{3}-\frac{1}{2}\alpha^{2}}\pi_{0}(s_{2})}.$ The comparative statics for the prior follows intuitively. For example, a higher $\pi_{0}(s_{1})$ leads to greater cross price impact between securities $x_{1}$ and $x_{3}$ by causing the market maker to place more emphasis on the co-movement of the orders for $x_{1}$ and $x_{3}$ in his inference.
@@ 4.2 Sufficient Statistics
The market maker adjusts for the noise trading intensities in making cross-market inferences. Higher noise levels reduce the informativeness of his sufficient statistic $(\sum\limits_{i}\frac{\widetilde{W}(x_{i}|s_{k})\omega_{i}}{\sigma_{i}^{2}})_{s_{k}\in S}$. In the limit where $\sigma_{i}\!\rightarrow\!\infty$ for each $i$, this statistic reduces to the trivial zero statistic, and his posterior (11) reduces to the prior $\pi_{0}$, i.e., he infers no information from order flow.
@@ 5.3 No-Arbitrage
No-arbitrage is a general necessary condition for the existence of equilibrium. We now formulate this condition for our setting. Here, once again, the projection coefficient profile $\Pi_{insider}$ between the insider’s trades and the market maker’s belief naturally occurs. Given market maker belief $\widetilde{W}$, we define the corresponding zero price impact portfolios as portfolios $W$ for which $\Pi_{insider}(W,\,\cdot\,;\widetilde{W})=0$. Intuitively, such a portfolio causes zero price impact because it is orthogonal to the market maker’s belief about the insider’s trading strategy. These zero price impact portfolios must yield a zero payoff for the insider; otherwise, they would present arbitrage opportunities. This is formally stated in Theorem 5.4.
@@
For any given market maker belief, an optimal portfolio for the insider can exist only if the corresponding zero price impact portfolios yield zero payoff for the insider. Otherwise, the insider could obtain unbounded utility.
@@ 6.1 Canonical Game
The zero payoff portfolios are those orthogonal to the linear span of $\{\eta(\,\cdot\,|s_{k})\}_{k}$—any portfolio $W$ satisfying $\sum\limits_{i}W(x_{i})\eta(x_{i}|s_{k})=0$ for all $s_{k}$ yields zero payoff for the insider. Thus, the insider’s payoff-relevant portfolios lie within the linear span of $\{\eta(\,\cdot\,|s_{k})\}_{k}$. We state this result as Proposition 6.1(i). On the other hand, the market maker’s noise-adjusted equilibrium beliefs must contain this span; otherwise, zero price impact allows arbitrage by the insider. This is Proposition 6.1(ii).
@@ 7.1 Informed Demand
Informed $\text{{\em Informed Demand}}\,\begin{cases}\text{{\em Single-Asset} :}&\frac{\sigma_{\varepsilon}}{\sigma_{v}}\cdot(s-v_{0})\\[2.58334pt] \text{{\em AD Securities} :}&{\bf L}^{-1}\cdot\alpha^{*}{\bf Q}\end{cases}.$
  • Information Intensity ($\frac{\sigma_{v}}{\sigma_{\varepsilon}}$ vs. ${\bf L}$): Both $\frac{\sigma_{v}}{\sigma_{\varepsilon}}$ and ${\bf L}$ quantify the insider’s information intensity in their respective settings. In the single-asset setting, a larger $\frac{\sigma_{v}}{\sigma_{\varepsilon}}$ means the market maker can more effectively infer the variation in the asset payoff from the variation in order flow, thus reducing informed demand. For contingent claims, a larger ${\bf L}$ means the market maker can better infer the variations in the payoff distributions across signals from the variations in cross-market order flow, which reduces informed demand.
@@
The following examples illustrate the informed demand for trading on the first, second, and third moments of an underlying asset payoff. The informed demands mirror widely used option strategies, demonstrating that they can indeed be explained in an equilibrium framework. Although we can easily replicate actual option strategies exactly by tweaking the distributions, we adhere to payoff distributions from the standard normal family, resulting in more stylized illustrations. Since our model imposes no parametric restrictions, in similar fashion we can explain a broad range of common trading strategies by adapting the model specifications. This provides a comprehensive theoretical framework that encompasses observed options trading practices, while also delivering actionable insights for designing option strategies. In Examples 7.1, 7.2, and 7.3, the signal is binary ($S=\{s_{1},s_{2}\}$), and all payoff distributions are suitably discretized.
@@
(a)/(b): high expected payoff signal/bull spread; (c)/(d): low expected payoff signal/bear spread.
@@
(a)/(b): high volatility signal/straddle; (c)/(d): low volatility signal/butterfly.
Price Impact Three states are indicated: $x_{1}$, $x_{2}$, and $x_{3}$.
@@
$\bullet\,x_{2}$: zero price impact on all securities because its payoff is constant (has zero variation) across signals. Thus, the informed demand for $x_{2}$ must be zero. See (b) and (d).
@@
The left column of Figure 4 shows the possible payoff distributions, with the insider’s private signal shaded. The right column of Figure 4 shows the informed demand corresponding to each signal. The informed demands align with common option strategies for volatility trading (see Hull (2015)). When the insider anticipates high volatility, his option strategy in Figure 4b follows a straddle or long iron condor pattern. When he anticipates low volatility, his option strategy in Figure 4d follows a butterfly pattern.
@@ 7.2 Price Impact
    Informed Demand Scale $\alpha^{*}$: As the informed demand (33) scales up by $\alpha^{*}$, the price impact responds one-to-one.
@@
    Noise Intensity Adjustment $\frac{1}{\sqrt{\sigma_{i}\sigma_{j}}}$: Higher noise trading intensity in either market reduces cross price impact.
@@
    Payoff Covariance $\mathbb{E}\left[{\operatorname{\mathbb{Cov}}}\left(\eta(x_{i}|\,\cdot\,),\eta(x_{j}|\,\cdot\,)\big{|}\omega\right)\right]$: The (expected) covariance of securities $x_{i}$ and $x_{j}$ payoffs under the market maker’s equilibrium posterior.
@@ 9 Conclusion
In this paper, we unify the foundational elements of Arrow and Debreu (1954) and Kyle (1985) to develop a general model of joint price discovery across contingent claim markets. By connecting observed market practices with theoretical insights, our results bridge a longstanding gap between theory and empirics. Beyond its immediate empirical implications, our basic model offers a versatile platform for further extensions. For example, in a setting with risk-averse agents, where the contingent-claim prices aggregate risk as well as information, one can explore how information asymmetry distorts risk-sharing incentives across derivative markets. Extending our model to a dynamic environment provides a framework to study the temporal evolution of information spillovers across markets. These present pathways for future research.
@@ A.1 Model Assumptions
In other words, $\mathbb{P}_{\!\scriptscriptstyle 0}$ specifies the probability law of noise-only order flow $\omega$ when insider demand is zero across markets.
@@
The marginal utility functional $\partial\!\kern 0.80002ptJ(\,\cdot\,;W)\colon\mathbb{R}^{\infty}\rightarrow\mathbb{R}$ is the Gâteaux-derivative of the expected utility functional $J(\,\cdot\,|s_{k};\widetilde{W})\colon\mathbb{R}^{\infty}\rightarrow{\mathbb{R}}$, as defined in Equation (14), at $W$. (Both $J(\,\cdot\,|s_{k};\widetilde{W})$ and its Gâteaux-derivative $\partial\!\kern 0.80002ptJ(\,\cdot\,;W)$ are conditional on $s_{k}$ and parameterized by $\widetilde{W}$.)
$J(W|s_{k};\widetilde{W})=J_{p}(W)-J_{c}(W)$ is the difference between the payoff functional
$J_{p}(W)=\sum\limits_{i}W(x_{i})\eta(x_{i}|s_{k})$
and the cost functional
$J_{c}(W)=\sum_{i}W(x_{i})\overline{P}(x_{i},W;\widetilde{W}).$